# Scalable Data-Driven Basis Selection for Linear Machine Learning Interatomic Potentials

### v1

@@ III Results
TABLE I: Summary of solvers used for ACE model construction.
@@
RRQR Rank-revealing QR decomposition with random matrix sketching; removes sensitive subspaces based on $rtol$, which is
related to $\lambda$ (Eq. II-A[13]. More performant than standard QR for large datasets.
@@
BLR Bayesian linear regression; selects $\lambda$ (Eq. II-A) via Bayesian evidence maximization [22, 34]. More robust than RRQR but
more computationally intensive. Best for small datasets and uncertainty quantification.
@@
ARD Empirical Bayes method with hierarchical sparsity constraints to prune basis functions [33].
@@
ASP Homotopy solver for LASSO; refines active basis set along a continuous path [27]. Scalable and well-suited for large
datasets requiring structured sparsity selection.
@@
OMP Greedy sparse regression method, selecting the most relevant basis functions iteratively [27]. More computationally
efficient than ASP.
@@
A summary of all the solvers used within this section and their most appropriate use cases is shown in Table I. This table provides a concise overview of each solver’s methodology, along with guidance on when it is most applicable for different problem settings. For each set of results we compare sparse solvers to a reference dense solver employed in previous tests in [52].
III-A Limited diversity metal datasets
@@
We evaluate the performance of the ASP and OMP solvers with default parameters on several low-diversity datasets, which arise for example in automated workflows that require special purpose potentials for a limited range of tasks, in particular (hyper-)active learning. We employ the single-element benchmarks [55] including six datasets for, respectively, Li, Mo, Ni, Cu, Si, and Ge. These elements cover a range of chemical properties (main group metals, transition metals, and semiconductors), crystal structures (bcc, fcc, and diamond), and bonding types (metallic and covalent). Each dataset comprises the element’s ground-state crystal structure, strained configurations with strains ranging from $-10\%$ to $10\%$, slab structures with Miller indices up to three, and NVT ab initio molecular dynamics simulations of bulk supercells, both with and without a single vacancy. We refer to [55] for further details. While the datasets provide a large number of training structures, they exhibit limited diversity compared to more recent MLIPs training sets.
@@
Our starting basis size was approximately 5000 (corresponding to a $D_{\rm tot}$ of 25 and $N_{\rm max}$ of 3). We then allowed OMP, ASP, and the ARD solver from scikit-learn [39] to select the most relevant basis functions. We terminated ASP and OMP when the number of active elements hit 300 for the small model and 1000 for the large model. We also obtained the ARD solutions with desired sparsity by modifying $\bm{\alpha}$. The resulting MAE errors of the potentials trained are compared to the previous results reported in [52]. As can be seen from the Tables II(a)II(b), apart from Li where the RRQR solver outperforms OMP and ASP, which suggests that fine-tuning of the model parameters could be beneficial, the OMP and ASP solvers outperform all other elements’ best checkpoint reported in [52].
@@
TABLE II: Mean absolute test errors in predicted energies (meV) and forces (eV/Å) for small ACE models ($\approx 300$ basis functions) and large ACE models ($\approx 1000$ basis functions) using different solvers, compared to the two best-performing MLIPs reported in [55].
@@
Figures 2 and 3 visualize the basis functions selected by the ASP solver for, respectively, the two-body and three-body interactions. The patterns of selected basis functions are non-intuitive and deviate significantly from any common a priori basis selection mechanism such as total degree or hyperbolic cross selections. This highlights the significant potential benefits of applying data-driven basis selection methods.
@@
Figure 1: Energy MAE vs basis size for selected limited diversity datasets (cf. III-A), comparing three sparse least squares solvers (ARD, ASP, OMP) with a direct regularized least squares approach (RRQR).
@@
Figure 2: Visualization of the basis functions selected for two-body interactions in the Mo dataset (cf.  III-A). The figure illustrates the gradual selection process of ASP and non-sparse ACE solvers as the number of active basis functions increases. The ASP-selected basis functions (pink) show a distinct, data-driven selection pattern compared to the full ACE basis (black), demonstrating the benefits of data-driven basis selection.
@@
Figure 3: Visualization of the basis functions selected for three-body interactions in the Mo dataset (cf.  III-A). The figure illustrates the gradual selection process of ASP and non-sparse ACE solvers as the number of active basis functions increases. The selection is visualized in 3D, with colors indicating the distance from the origin. Similarly as with 2-correlation, ASP selects features without any a priori predictable pattern.
III-B Silicon (PRX, 2018)
@@
We employed the ASP and OMP solvers to fit a linear ACE potential to the silicon dataset presented in [4], originally developed for training a Gaussian approximation potential (GAP). This dataset is comprehensive, encompassing a diverse range of configurations, including various bulk crystal structures (e.g., diamond, hcp, and fcc), amorphous phases, and liquid molecular dynamics (MD) snapshots. Its goal is to capture a wide range of the elemental silicon energy landscape. Accurate fits to this dataset provide (near-)general-purpose potentials for elemental silicon.
Figure 4: Energy MAE vs. basis size for the Silicon dataset [4], comparing OMP, BLR, and ASP (cf. III-B). In the mid-to-high basis size regime, ASP and OMP achieve similar or better accuracy than BLR while using significantly fewer basis functions.
@@
Four models were evaluated: the GAP model and three ACE models trained with default parameters. The ACE model trained using BLR employed basis functions with a maximum correlation order of $N_{\text{max}}=3$, a total polynomial degree of $D_{\text{tot}}=20$, and a cutoff radius of $r_{c}=6\,\text{\AA}$, resulting in a basis size of 5,434.
Additionally, two more ACE models were trained using ASP and OMP, both with the same maximum correlation order of 3 but a higher total degree of $D_{\text{tot}}=23$, yielding a larger initial basis size of 13,695. In both cases, training was stopped once 5,000 active basis functions had been selected.
An algebraic smoothness prior $\Gamma$ with $p=5$ was employed in all ACE models. Figure 4 illustrates the comparison of energy Mean Absolute Error (MAE) between the ASP, OMP, and BLR solver, across varying basis sizes. In the small basis regime, the BLR, OMP, and ASP solvers exhibit comparable performance in terms of test error. As the basis size increases, the ASP and OMP solver consistently achieve moderately lower test errors compared to BLR, which indicates that selecting an effective subset of basis functions leads to better generalization and less risk of overfitting, even in the current setting with a highly diverse dataset.
@@
TABLE III: Comparison of point defects, elastic properties, and surface energies for silicon using DFT, GAP, BLR, ASP, and OMP models (cf. III-B). Models were fitted to the dataset in [4], with reference values obtained from CASTEP [20] density functional theory (DFT) calculations.
@@
A key property of a robust interatomic potential is its out-of-distribution generalization performance. To evaluate this, the benchmark suite of property predictions from [4] was performed. It includes bulk diamond elastic constants, vacancy formation energies, surface formation energies for the (100), (110), and (111) surfaces, as well as formation energies for hexagonal, dumbbell, and tetragonal point defects in bulk diamond. The reference values for these properties were obtained via CASTEP [20] density functional theory (DFT) calculations.
For comparison, we included results from the GAP potential trained in [4] and the ACE BLR potential trained in [52]. The computed values of these properties are summarized in Table III. These results demonstrate that the ASP and OMP solvers achieve accuracy comparable to the GAP potential and the ACE model trained using BLR, even when training is stopped at 2,500 basis functions—less than $50\%$ of the 5,434 basis functions used by the BLR-trained model.
Figure 5: Percentage error relative to the computed values in Table III for various silicon properties using GAP, BLR, and ASP (cf. III-B). The plot illustrates how the relative error of ASP decreases as the number of active basis functions increases, demonstrating a clear trend of improved accuracy.
@@ III-C Water
We utilize OMP and ASP to fit an interatomic potential to a dataset containing 1593 liquid water configurations [18]. In this case, we can test the selection of how basis functions acting between different elements are selected. The model employed default parameters, including basis functions up to a correlation order of $N_{\text{max}}=3$, a maximum polynomial degree of $D_{\text{tot}}=15$, and a cutoff radius of $r_{\text{c}}=5.5$. An algebraic smoothness prior $\Gamma$ with $p=4$ was employed in all ACE models.
Figure 6 presents the learning curve for ACE models trained on the water dataset, comparing the performance of OMP, ASP, and BLR solvers. Both OMP and ASP exhibit comparable mean absolute errors (MAE) while significantly outperforming BLR. In this case, OMP achieved slightly better accuracy than ASP while being more computationally efficient, hence we present all further results using only the OMP solver.
@@
In the following numerical test, ASP and OMP models were initialized with a basis size of approximately 19,000, and training was terminated once the number of active basis functions reached 3,000, 9,000, and 12,000. Table IV compares the results obtained using ASP, OMP, and BLR with Cartesian Atomic Cluster Expansion (CACE) [19], which is a recently proposed framework that formulates atomic density expansions directly in Cartesian coordinates, bypassing the need for spherical harmonics. CACE $T=0$ relies on local atomic features (similar to linear ACE), while $T=1$ incorporates message passing to capture non-local interactions, improving accuracy at a higher computational cost.
@@
Notably, the ACE models fitted using OMP require significantly fewer parameters than other models in Table IV. Specifically, the OMP solver achieves lower energy error than the CACE ($T=0$) model while using less than 50% of the parameters. This is despite being a linear model, whereas CACE employs nonlinear tensor decompositions and a learnable radial basis. The larger force error suggests that tuning the regression weights may be beneficial, a process that should arguably also be automated in future work.
@@
Figure 6: Energy MAE vs. basis size for the Water dataset [18], comparing OMP, BLR, and ASP (cf. III-C). ASP and OMP Significantly outperform BLR while using significantly fewer basis functions.
@@
Figure 7: Illustration of ASP and other (non-sparse) ACE solvers’ gradual selection of three-body basis functions for the water dataset [18], with O as the center atom (cf. III-C). Each row corresponds to a different interaction type: O-H-H, O-O-O, and O-O-H. The results highlight the greater importance of resolving O-H interactions for accurately modeling the potential energy surface. This aligns with chemical intuition, as these bonds dominate water’s structural and energetic properties due to their role in hydrogen bonding and intermolecular interactions.
TABLE IV: Test Mean Absolute Errors (MAE) for energy (E) per water molecule (in meV/H2O) and force (in meV/Å) for models trained on a liquid water dataset from Ref. [18]. The cutoff radius $r_{\mathrm{cut}}$ of the atomic environment and the number of message passing layers $T$ are listed. Despite being a linear model, the sparse solver achieves comparable accuracy with significantly fewer parameters, demonstrating the effectiveness of data-driven basis selection.
Model $r_{\mathrm{cut}}$ [Å] Energy MAE (meV/atom) Force MAE (meV/Å) # Parameters
CACE [19] ($T=1$) 5.5 0.89 27 69320
CACE [19] ($T=0$) 5.5 1.43 42 24572
ACE (BLR) [52] 5.5 1.86 76 19501
ACE (OMP) 5.5 2.02 88 3000
ACE (OMP) 5.5 1.43 63 9000
ACE (OMP) 5.5 1.35 59 12000
IV Conclusion

### v2

@@ 3 Results
Table 1: Summary of solvers used for ACE model construction.
@@
RRQR Rank‐revealing QR decomposition with random matrix sketching; removes sensitive subspaces based on the tolerance parameter $r_{\mathrm{tol}}$, which is related to $\lambda$ (Eq. 2.1[14]. More performant than standard QR for large datasets.
@@
BLR Bayesian linear regression that selects $\lambda$ via Bayesian evidence maximization [24, 41]. More robust than RRQR but computationally intensive; best suited for small datasets and uncertainty quantification.
@@
ARD Empirical Bayes method with hierarchical sparsity priors to prune basis functions automatically [40].
@@
ASP Homotopy solver for the LASSO problem; refines the active basis set along a continuous path [30]. Scalable and well‐suited for large datasets requiring structured sparsity.
@@
OMP Greedy sparse regression algorithm that selects the most relevant basis functions iteratively [30]. More computationally efficient than ASP.
@@
A summary of all the solvers used within this section and their most appropriate use cases is shown in Table 1. This table provides a concise overview of each solver’s methodology, along with guidance on when it is most applicable for different problem settings. For each set of results we compare sparse solvers to a reference dense solver employed in previous tests in [62].
3.1 Limited diversity material datasets
@@
We evaluate the performance of the ASP and OMP solvers with default parameters on several low-diversity datasets, which arise for example in automated workflows that require special purpose potentials for a limited range of tasks, in particular (hyper-)active learning. We employ the single-element benchmarks [65] including six datasets for, respectively, Li, Mo, Ni, Cu, Si, and Ge. These elements cover a range of chemical properties (main group metals, transition metals, and semiconductors), crystal structures (bcc, fcc, and diamond), and bonding types (metallic and covalent). Each dataset comprises the element’s ground-state crystal structure, strained configurations with strains ranging from $10\%$ to $10\%$, slab structures with Miller indices up to three, and NVT ab initio molecular dynamics simulations of bulk supercells, both with and without a single vacancy. We refer to [65] for further details. While the datasets provide a large number of training structures, they exhibit limited diversity compared to more recent MLIPs training sets.
@@
For this benchmark, the training and test sets are predefined. To construct the validation set, we randomly shuffle the training data and split it into 85% training and 15% validation subsets. Our starting basis size was 5292 (corresponding to a $D_{\rm tot}$ of 25 and a body order of 3). We then allowed OMP, ASP, and the ARD solver from scikit-learn [47] to select the most relevant basis functions. We terminated ASP and OMP when the number of active elements hit 300 for the small model and 1000 for the large model. We also obtained the ARD solutions with desired sparsity by modifying $\boldsymbol{\alpha}$. The resulting MAE errors of the potentials trained are compared to the previous results reported in [62]. As can be seen from the Tables 2(a)2(b), apart from Li where the RRQR solver outperforms OMP and ASP, which suggests that fine-tuning of the model parameters could be beneficial, the OMP and ASP solvers outperform all other elements’ best checkpoint reported in [62].
@@
Table 2: Mean absolute test errors in predicted energies (meV) and forces (eV/Å) for small ACE models ($\approx 300$ basis functions) and large ACE models ($\approx 1000$ basis functions) using different solvers, compared with the two best-performing MLIPs reported in [65].
@@
Figures 2 and 3 visualize the basis functions selected by the ASP solver for, respectively, the two-body and three-body interactions. In Figure 2, the black squares correspond to basis functions selected by any non-sparse solver at a given basis size, while the pink squares indicate the basis functions chosen by the ASP solver. In Figure 3, the dots correspond to the basis functions chosen by ASP, separated according to their distance from the origin for visualization purposes.
The ASP solver tends to select a larger proportion of three-body basis functions compared to two-body ones. The patterns of selected basis functions are non-intuitive and deviate significantly from any common a priori basis selection mechanism such as total degree or hyperbolic cross selections. A speculative explanation is that two-body features can be approximately represented in terms of three-body features, hence the sparse basis selection reduces the resulting ill-conditioning. Regardless of the explanation for the selected basis, this result highlights the significant potential benefits of applying data-driven basis selection methods.
@@
Figure 1: Energy MAE vs basis size for selected limited diversity datasets (cf. 3.1), comparing three sparse least squares solvers (ARD, ASP, OMP) with a direct regularized least squares approach (RRQR).
@@
Figure 2: Visualization of the basis functions selected for two-body interactions in the Mo dataset (cf.  3.1). The figure illustrates the gradual selection process of ASP and non-sparse ACE solvers as the number of active basis functions increases. The ASP-selected basis functions (pink) show a distinct, data-driven selection pattern compared to the full ACE basis (black), demonstrating the benefits of data-driven basis selection.
@@
Figure 3: Visualization of the basis functions selected for three-body interactions in the Mo dataset (cf.  3.1). The figure illustrates the gradual selection process of ASP and non-sparse ACE solvers as the number of active basis functions increases. The selection is visualized in 3D, with colors indicating the distance from the origin. Similarly as with 2-correlation, ASP selects features without any a priori predictable pattern.
We performed an additional test on the limited-diversity elements dataset, starting the OMP solver with varying initial basis sizes to examine their effect on model accuracy when the solvers were terminated at approximately 1000 basis functions. The results are shown in Table 3. In some cases, we observe a slight decrease in energy error with increasing initial basis size, while the force error remains stable. However, these changes are small, suggesting that the starting basis size does not have a significant impact on accuracy. This indicates that users can start the solvers with the largest basis size suitable for their application. In our own tests we generally use a starting basis approximately twice as large as the target basis size.
Table 3: Energy and force mean absolute errors (MAE) for different numbers of starting basis functions for the Mo, Ni, and Si datasets. All models select 1000 features. Increasing the starting basis size gives consistent moderate improvements in fit accuracy, suggesting that it should be chosen primarily based on affordable computational cost for parameter estimation.
Element # Starting basis Energy MAE (meV/atom) Force MAE (meV/Å)
Mo 1429 2.756 0.092
2275 1.952 0.091
3515 2.101 0.090
5292 1.933 0.089
7777 2.242 0.088
Si 1429 2.637 0.065
2275 2.403 0.063
3515 1.775 0.061
5292 1.859 0.062
7777 1.558 0.062
Ni 1429 0.297 0.014
2275 0.217 0.014
3515 0.229 0.014
5292 0.241 0.013
7777 0.276 0.013
3.1.1 Comparison with Other Path-Tracing Solvers
To explain our choice of the ASP and OMP solvers, we compared them against two available software packages that generate LASSO paths. The results are summarized in Table 4. As shown, the MAEs obtained from potentials fitted with LARS.jl [36] and Lasso.jl [35] are significantly higher than those achieved with ASP and OMP.
Table 4: Energy and force mean absolute errors (MAE) obtained using different sparse regression packages on the silicon and molybdenum datasets [65]. The ASP and OMP solvers are from ActiveSetPursuit.jl [57], the LARS solver is from LARS.jl [36], and the LASSO path solver is from Lasso.jl [35]. All reported models were trained using an initial basis of 5292 functions and selected approximately 1000 active features.
(a) Mean absolute test errors in predicted energies (meV).
Element Small model [meV] Large model [meV]
ASP OMP LARS LASSO ASP OMP LARS LASSO
Mo 2.261 2.123 20.627 21.066 2.138 1.933 12.892 13.787
Si 2.444 2.296 19.916 20.111 1.981 1.859 12.970 13.326
(b) Mean absolute test errors in predicted forces (eV/Å).
Element Small model [eV/Å] Large model [eV/Å]
ASP OMP LARS LASSO ASP OMP LARS LASSO
Mo 0.104 0.098 0.247 0.254 0.087 0.086 0.152 0.159
Si 0.074 0.072 0.284 0.290 0.061 0.059 0.168 0.177
3.2 Silicon (PRX, 2018)
@@
We employed the ASP and OMP solvers to fit a linear ACE potential to the silicon dataset presented in [4], originally developed for training a Gaussian approximation potential (GAP). This dataset is more diverse and distinct from the Si dataset used in Section 3.1. This dataset is comprehensive, encompassing a diverse range of configurations, including various bulk crystal structures (e.g., diamond, hcp, and fcc), amorphous phases, and liquid molecular dynamics (MD) snapshots. Its goal is to capture a wide range of the elemental silicon energy landscape. Accurate fits to this dataset provide (near-)general-purpose potentials for elemental silicon. For this benchmark, the data are randomly divided into 70% training, 15% validation, and 15% test sets. The split respects training structure categories, such as different phases or defects.
@@
Four models were evaluated: the GAP model and three ACE models trained with default parameters. The ACE model trained using BLR employed basis functions with a maximum correlation order of $N_{\text{max}}=3$, a total polynomial degree of $D_{\text{tot}}=20$, and a cutoff radius of $r_{c}=6\,\text{\AA }$, resulting in a basis size of 5,456.
Additionally, two more ACE models were trained using ASP and OMP, both with the same maximum correlation order of 3 but a higher total degree of $D_{\text{tot}}=23$, yielding a larger initial basis size of 13,695. In both cases, training was stopped once 5,000 active basis functions had been selected. An algebraic smoothness prior $\Gamma$ with $p=5$ was employed in all ACE models. A key property of a robust interatomic potential is its out-of-distribution generalization performance. To evaluate this, the benchmark suite of property predictions from [4] was performed. It includes bulk diamond elastic constants, vacancy formation energies, surface formation energies for the (100), (110), and (111) surfaces, as well as formation energies for hexagonal, dumbbell, and tetragonal point defects in bulk diamond. The reference values for these properties were obtained via CASTEP [21] density functional theory (DFT) calculations.
For comparison, we included results from the GAP potential trained in [4] and the ACE BLR potential trained in [62]. The computed values of these properties are summarized in Table 5.
@@
Table 5: Comparison of point defects, elastic properties, and surface energies for silicon using DFT, GAP, BLR, ASP, and OMP models (cf. 3.2). Models were fitted to the dataset in [4], with reference values obtained from CASTEP [21] density functional theory (DFT) calculations.
@@
These results demonstrate that the ASP and OMP solvers achieve accuracy comparable to the GAP potential and the ACE model trained using BLR, even when training is stopped at 2,500 basis functions—less than $50\%$ of the 5,456 basis functions used by the BLR-trained model.
Figure 4: Percentage error relative to the computed values in Table 5 for various silicon properties using GAP, BLR, and ASP (cf. 3.2). The plot illustrates how the relative error of ASP decreases as the number of active basis functions increases, demonstrating a clear trend of improved accuracy.
@@
Figure 5 compares OMP, ASP, and BLR fits for silicon. For OMP and ASP, the solvers are initialized with a basis of 18,353 functions, and the test error is evaluated along the solution path up to 10,000 active basis functions. Results are averaged over 7 independent train–test splits, and shaded regions indicate 95% confidence intervals for the mean test error. In the small-basis regime, BLR and ASP exhibit comparable performance in terms of test error. As the basis size increases, the ASP and OMP solver consistently achieve moderately lower test errors compared to BLR, which indicates that selecting an effective subset of basis functions leads to better generalization and less risk of overfitting, even in the current setting with a highly diverse dataset.
Figure 5: Energy MAE as a function of basis size for the silicon dataset [4]. Mean test errors are averaged over 7 train–test splits, with shaded regions indicating 95% confidence intervals for the mean. ASP and OMP outperform BLR while using significantly fewer basis functions.
@@ 3.3 Water
We utilize OMP and ASP to fit an interatomic potential to a dataset containing 1593 liquid water configurations [19]. In this case, we can test the selection of how basis functions acting between different elements are selected. The model employed default parameters, including basis functions up to a correlation order of $N_{\text{max}}=3$, a maximum polynomial degree of $D_{\text{tot}}=15$, and a cutoff radius of $r_{\text{c}}=5.5$ Å. An algebraic smoothness prior $\Gamma$ with $p=4$ was employed in all ACE models.
We compared the performance of the ASP, OMP, and BLR solvers using 7 independent random splits of the dataset, each partitioned into 70% training, 15% validation, and 15% test sets. Figure 7 shows the learning curves for ACE models trained on the water dataset, starting from an initial basis of 19,501 functions, with the test error evaluated along the solution path up to 10,000 active basis functions. Results are averaged over the 7 splits, and shaded regions indicate 95% confidence intervals for the mean. OMP and ASP achieve comparable MAE and consistently outperform BLR. As OMP offers similar predictive accuracy while being more computationally efficient, we present all further results using only the OMP solver.
We next compared the OMP solver with Cartesian Atomic Cluster Expansion (CACE) [20], which is a recently proposed framework that formulates atomic density expansions directly in Cartesian coordinates, bypassing the need for spherical harmonics.
In the following numerical test, the OMP model was initialized with a basis of 19,501 functions, and training was terminated once the number of active basis functions reached 3,000, 9,000, and 12,000.
For a fair comparison, the dataset [19] was partitioned into 1,354 training configurations and 239 test configurations, which were used to train and evaluate the OMP, BLR, and CACE models. A validation set comprising 15% of the training data was used for model selection.
@@
Table 6 compares the results obtained using OMP and BLR with CACE. CACE $T=0$ relies on local atomic features (similar to linear ACE), while $T=1$ incorporates message passing to capture non-local interactions, improving accuracy at a higher computational cost.
@@
Notably, the ACE models fitted using OMP require significantly fewer parameters than other models in Table 6. Specifically, the OMP solver achieves lower energy error than the CACE ($T=0$) model while using less than 50% of the parameters. This is despite being a linear model, whereas CACE employs nonlinear tensor decompositions and a learnable radial basis. The larger force error suggests that tuning the regression weights may be beneficial, a process that should arguably also be automated in future work.
Table 6: Test Mean Absolute Errors (MAE) for energy (E) per water molecule (in meV/H2O) and force (in meV/Å) for models trained on a liquid water dataset from Ref. [19]. The number of message passing layers $T$ for the CACE models are listed. All models used a cutoff radius of $r_{\text{cut}}=5.5\text{\AA }$. Despite being a linear model, the sparse solver achieves comparable accuracy with significantly fewer parameters, demonstrating the effectiveness of data-driven basis selection.
Model Energy MAE (meV/atom) Force MAE (meV/Å) # Parameters
CACE [20] ($T=1$) 0.89 27 69320
CACE [20] ($T=0$) 1.42 42 24572
ACE (BLR) [62] 1.86 76 19501
ACE (OMP) 1.58 71 6000
ACE (OMP) 1.44 65 9000
ACE (OMP) 1.36 62 12000
To test the stability of trained potentials, we ran NVT MD simulations at 300 K with a timestep of 1 fs for a total simulation time of 300 ps. From these simulations, we computed the diffusivities reported in Table 7.
Table 7: Diffusivities obtained from NVT MD simulations at 300 K using different potentials, compared with the DFT reference value. Overall, diffusivity errors follow the same trend as force errors, with CACE ($T=0$) being the only exception.
Model Diffusivity (Å2/ps) Relative Diffusivity error
ACE BLR 0.120 0.47
ACE OMP (6000) 0.101 0.55
ACE OMP (9000) 0.118 0.47
ACE OMP (12000) 0.129 0.42
CACE ($T=0$) 0.168 0.25
CACE ($T=1$) 0.174 0.23
DFT [43] 0.226 -
As the number of active basis functions increases, the computed diffusivity approaches the DFT value. The diffusivity errors follow essentially the same trend as the force errors, with models that achieve lower force MAEs reported in Table 6, generally producing more accurate predictions. The diffusivity accuracy of linear ACE models is generally poor compared with the nonlinear CACE models, which is consistent with previous studies [62]. Most importantly for our purposes is that all potentials are sufficiently MD stable to be able to predict diffusivity and that the OMP potential matches or outperforms the dense BLR ACE potential in accuracy with significantly fewer active basis functions.
@@
Figure 6: Illustration of ASP and other (non-sparse) ACE solvers’ gradual selection of three-body basis functions for the water dataset [19], with O as the center atom (cf. 3.3). Each row corresponds to a different interaction type: O-H-H, O-O-O, and O-O-H. The results highlight the greater importance of resolving O-H interactions for accurately modeling the potential energy surface. This aligns with chemical intuition, as these bonds dominate water’s structural and energetic properties due to their role in hydrogen bonding and intermolecular interactions.
Figure 7: Energy MAE vs. basis size for the water dataset [19], comparing OMP, BLR, and ASP (cf. 3.3). ASP and OMP outperform BLR while using significantly fewer basis functions. Solid lines indicate the mean error over 7 train–test splits, with shaded regions indicating 95% confidence intervals for the mean.
3.4 Effects of correlation order
To examine the role of the physics encoded in the descriptor set, we performed additional sparse OMP fits using correlation orders $N_{\max}=3,4,5$ for both silicon and water. For each case, the polynomial degree was chosen such that the total number of candidate basis functions was comparable (approximately $2\times 10^{4}\pm 2\times 10^{3}$). For each correlation order, we trained 7 models using random train-test splits; with shaded regions indicating 95% confidence intervals for the mean. The resulting learning curves, shown in Fig. 8, demonstrate that our method is robust as the underlying dimensionality of the model increases. Comparable accuracy is achieved across correlation orders once an appropriate sparse subset of features is selected.
Figure 8: Energy MAE as a function of basis size for OMP models trained with different correlation orders $N_{\max}=3,4,5$. (a) shows results for the Silicon dataset [4], and (b) for the water dataset [19]. In both systems, increasing the correlation order does not yield a consistent improvement in test-set accuracy at fixed sparsity. Solid lines indicate the mean error over 7 train–test splits, with shaded regions indicating 95% confidence intervals for the mean.
On the other hand, increasing the correlation order does not lead to a reduction in the error at fixed sparsity for either system. This behavior is in fact unsurprising: As shown in Figs. 2 and 6, the sparse solvers tend to select high-degree basis functions at the lower correlation orders. When the correlation order is increased while keeping the total basis size fixed, these high-degree basis functions are no longer available in the candidate pool. As a result, increasing the correlation order does not improve accuracy at fixed sparsity.
4 Conclusion